# Lumos-1: On Autoregressive Video Generation from a Unified Model Perspective

### v1

@@ Appendix A Detailed AR-DF Inference Algorithm
In this subsection, we present the detailed version of the AR-DF inference algorithm. After the text prompt $\bm{X}_{p}$ is encoded once and its KV pairs are cached (Lines 1–3), we iterate over the $T$ latent frames (Line 6). Each frame is generated by running $N_{steps}$ iterations (Lines 9–22). During the frame generation process, the model predicts a token distribution, samples tokens only at positions that are still masked, gathers their confidences, and sets the scores of already-revealed tokens to +$\infty$+ ∞. After adding Gumbel noise, we re-mask the lowest-confidence tokens. As the generation proceeds, fewer tokens are re-masked as the running steps grow and the frame becomes sharper and clearer. The partially revealed frame is then added to the cache before the next frame is generated. Once all frames are finished, they are decoded to RGB space to form the final video.
@@
3:$\mathcal{C}\leftarrow G_{\phi}(\bm{X}_{p},AttnMask^{(p)})$triangleright Store cache for the prompt
@@
10:         $\bm{P}\leftarrow G_{\phi}\bigl{(}\bm{X}_{v}^{(t)},AttnMask^{(t)},\mathcal{C}\bigr{)}$triangleright Predicted token distribution for all tokens
[…]
@@
17:          $\alpha\leftarrow\cos(\frac{\pi}{2}\frac{n}{N_{steps}})$triangleright Cosine schedule factor in [0,1]
18:          $U\leftarrow\text{floor}(\alpha\times N_{f})$triangleright Number of tokens to mask in this iteration
[…]
@@
26:     $\mathcal{C}\leftarrow G_{\phi}(\widetilde{\bm{X}}_{v}^{(t)},AttnMask^{(t)},\mathcal{C})$triangleright Store cache for the $t$-th frame
@@
Figure A1: Detailed inference algorithm with AR-DF.
Appendix B More Architectural and Implementation Details
Model architectures and more training details. Tab. B1 presents the architectural details of three versions of Lumos-1 and their corresponding training details, including their training batch sizes and their learning rates. For 256p joint training, we use all 60M images and 10M videos. For 384p joint training, we curate a higher-quality subset from 256p training data, consisting of 8M images and 0.6M videos. Following the alternating strategy of MovieGen [49], we interleave image and video batches during joint training.
@@ Appendix B More Architectural and Implementation Details
Table B1: Model architectural and training details. Batch sizes (*/*) are for images and videos during joint training. Learning rates (*/*) are for 256p and 384p. The 0.5B version is only for fast ablation studies.
@@
Lumos-1 3B 3.6B 28 3072 24 128 768 / 96 288 / 48 1e-4 / 2.5e-5
Appendix C More Visualizations
@@
Qualitative visual comparison on text-to-image generation. We compare Lumos-1 with popular text-to-image generation methods in figure C1. All models are configured to their default inference settings and use the same version of detailed prompts to ensure quality generation. We can observe that: 1) Although EMU3 is significantly larger than Lumos-1, Lumos-1 generates images that reproduces fine prompt details more faithfully, such as the dappled shadows on the bench (example 1) and the correct “STOP” text and the yellow color on the road sign (example 3). EMU-3 sometimes drops or warps these details and exhibits occasional anatomical distortions, such as the distorted zebra (example 5) and malformed fingers on the woman holding an apple (example 7). 2) Stable Diffusion 3 offers slightly crisper textures, which is expected from its continuous tokenizer, but it frequently ignores descriptive constraints, such as the dappled shadows (example 1) and gentle shadows (example 4), and the incorrect dress color (example 7). By contrast, Lumos-1 meets these constraints while maintaining a convincing overall composition. These examples illustrate that Lumos-1 delivers strong vision–language alignment compared with both diffusion and autoregressive baselines, while remaining competitive in visual quality.
@@
Figure C1: Visual comparison of Lumos-1 with EMU3 and Stabel Diffusion 3 (Medium) on the text-to-image task. We place the simple version of the prompts above the images for reference. Fragments of the detailed prompt are placed in parentheses, with the critical attributes highlighted in red. All images are generated using the detailed prompts for all models.
@@
More text-to-image visualizations. figure C2 presents more results generated by Lumos-1 using descriptive captions. Due to the unified codebook, we can train Lumos-1 using the original resolution of the visual data, enabling generation with different aspect ratios such as $4:7$, $7:4$, and $1:1$, although we do not have a significant amount of data with aspect ratio $4:7$. Even though Lumos-1 is trained from scratch, it transforms lengthy descriptions into complex scenes that are aligned faithfully with the prompts. This can be exemplified by the second row of images, where these images clearly reflect the prompts:
@@
Figure C2: More text-to-image results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The resolutions for three rows of images are $384\times 672$, $672\times 384$ and $512\times 512$.
More image-to-video and text-to-video visualizations. figure C3 presents more image-to-video results and figure C4 presents more text-to-video results generated by Lumos-1. The prompts span a wide semantic and dynamical range, containing 1) close-up animal/human scenes (example 4, 5 in figure C4), 2) fast human and vehicle motion (example 1, 4 in figure C3 and example 1, 2 in figure C4), 3) fluid dynamics (example 1, 5 in figure C3 and example 2 in figure C4), and 4) complex multi-object collective motion (example 3 in figure C3 and example 3 in figure C4). The results demonstrate that Lumos-1 can 1) handle videos of different aspect ratios and 2) generate single-subject and multi-subject motion that is temporally coherent and physically plausible.
Figure C3: More image-to-video results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The results encompass videos of resolution $672\times 384$ and $768\times 320$.
Figure C4: More text-to-video results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The resolution is $672\times 384$.
Appendix D Discussions
D.1 Limitations and Future Work Discussions
@@
D.2 Potential Societal Impacts and Safeguards

### v2

@@ Appendix C Detailed AR-DF Inference Algorithm
In this subsection, we present the detailed version of the AR-DF inference algorithm. After the text prompt $\bm{X}_{p}$ is encoded once and its KV pairs are cached (Lines 1–3), we iterate over the $T$ latent frames (Line 6). Each frame is generated by running $N_{steps}$ iterations (Lines 9–22). During the frame generation process, the model predicts a token distribution, samples tokens only at positions that are still masked, gathers their confidences, and sets the scores of already-revealed tokens to $\infty$. After adding Gumbel noise, we re-mask the lowest-confidence tokens. As the generation proceeds, fewer tokens are re-masked as the running steps grow and the frame becomes sharper and clearer. The partially revealed frame is then added to the cache before the next frame is generated. Once all frames are finished, they are decoded to RGB space to form the final video.
@@
3:   $\mathcal{C}\leftarrow G_{\phi}(\bm{X}_{p},AttnMask^{(p)})$ $\triangleright$ Store cache for the prompt
@@
10:    $\bm{P}\leftarrow G_{\phi}\bigl(\bm{X}_{v}^{(t)},AttnMask^{(t)},\mathcal{C}\bigr)$ $\triangleright$ Predicted token distribution for all tokens
11:    $\textit{sampled\_ids}\;\leftarrow\;\mathrm{multinomial}(\bm{P},\,1)$
12:    $\textit{sampled\_ids}\leftarrow\mathrm{where}\!\bigl(\bm{X}^{(t)}_{v}=[\text{MASK}],\;\textit{sampled\_ids},\;\bm{X}^{(t)}_{v}\bigr)$
13:    $\bm{Q}\leftarrow\mathrm{Gather}\bigl(\bm{P},\;\textit{sampled\_ids}\bigr)$ $\triangleright$ Select confidence score of every chosen token
14:    $\bm{Q}\leftarrow\mathrm{SetKnownInfinity}\!\bigl(\bm{Q},\;\bm{X}^{(t)}_{v}\neq[\text{MASK}]\bigr)$ $\triangleright$ Set $\infty$ where the token was not masked
@@
17:       $\alpha\leftarrow\cos(\frac{\pi}{2}\frac{n}{N_{steps}})$ $\triangleright$ Cosine schedule factor in [0,1]
18:       $U\leftarrow\text{floor}(\alpha\times N_{f})$ $\triangleright$ Number of tokens to mask in this iteration
19:       $\textit{low\_conf\_pos}\;\leftarrow\;\mathrm{SelectLowConfidence}\bigl(\bm{Q},\,U\bigr)$
20:       $\textit{sampled\_ids}[\textit{low\_conf\_pos}]\;\leftarrow\;[\text{MASK}]$
21:       $\bm{X}_{v}^{(t)}\;\leftarrow\;\textit{sampled\_ids}$
@@
26:     $\mathcal{C}\leftarrow G_{\phi}(\widetilde{\bm{X}}_{v}^{(t)},AttnMask^{(t)},\mathcal{C})$ $\triangleright$ Store cache for the $t$-th frame
@@
Figure C1: Detailed inference algorithm with AR-DF.
Appendix D More Architectural and Implementation Details
Model architectural details and more training details. For all of our models, we adopt the Adam [45] optimizer to optimize the model, set the weight decay to 0.1 and set $\beta_{1}$ and $\beta_{2}$ to $0.9$ and $0.95$. Tab.˜D1 presents the architectural details of three versions of Lumos-1 and their corresponding training details, including their training batch sizes and their learning rates. For 256p joint training, we use all 60M images and 10M videos. For 384p joint training, we curate a higher-quality subset from 256p training data, consisting of 8M images and 0.6M videos. Following the alternating strategy of MovieGen [53], we interleave image and video batches during joint training.
@@ Appendix D More Architectural and Implementation Details
GPU memory friendly implementation. By default, we leverage flash attention [14] to accelerate attention calculation and reduce memory overhead during training and inference of Lumos-1. Moreover, we observe significant amount of GPU memory consumption during the training of a model with a large codebook size. To address it, we eliminate the use of language-related loss (i.e., next-token prediction on texts), reducing the final logit matrix size to match only visual tokens. While text token embeddings, which map the text token index to token embeddings, remain trainable, this focuses the model on video generation. This loss can be added if we target a unified model that is capable of learning within the language modality. Finally, we observe substantially high GPU memory consumption during loss computation over 129kk token types, which easily leads to the out-of-memory issue. To solve this, we utilize chunked cross-entropy loss, which maintains full softmax accuracy by upcasting and calculating softmax logits one chunk of token sequences at a time. This approach significantly reduces peak memory usage. By default, we set the chunk size to 2,000.
Table D1: Model architectural and training details. Batch sizes (*/*) are for images and videos during joint training. Learning rates (*/*) are for 256p and 384p. The 0.5B version is only for fast ablation studies.
@@
Lumos-1 3B 3.6B 28 3072 24 128 768 / 96 288 / 48 1e-4 / 5.0e-5
Appendix E More Analysis and Ablation Studies
Figure E1: (a) Effect of the number of meta MM-RoPEs in $M$-RoPE; (b) Effect of the scaling factors in MM-RoPE. Both curves are plotted using the 0.5B model trained on videos.
E.1 More Ablation Studies
Effect of the number of meta MM-RoPEs in MM-RoPE. MM-RoPE slices the embedding channels into several meta groups. More groups mean that one kind of information (temporal, height or width) receives a wider and more comprehensive frequency spectrum for modeling. figure˜E1(a) plots the validation loss for a 0.5B model under four settings: 1) w/o distributed design: This setting uses the previous design, which allocates the first $\frac{2}{8}$ channels for temporal modeling and $\frac{3}{8}$ channels for height and width modeling, respectively. 2) #meta MM-RoPE=1: This setting equips a meta MM-RoPE of 64 channels, while maintaining the ratios ($2:3:3$) for modeling temporal, height and width information. This variant improves upon the previous one by enabling interleaved height and width channels, which increases the frequency spectrum range for the two spatial dimensions. 3) #meta MM-RoPE=2: This setting equips a meta MM-RoPE of 32 channels. This variant improves the frequency spectrum range for temporal, height and width information compared with the previous one. 4) #meta MM-RoPE=4: This setting is the default design that keeps the number of channels for every meta MM-RoPE minimal (16 channels). This means that the frequency spectrum range is most comprehensive for the temporal, height, or width dimension. These results confirm that widening the frequency spectrum for every dimension through increasing the number of meta MM-RoPEs substantially improves spatiotemporal modeling and overall training effectiveness.
Table E1: GenEval score across different aspect ratios using 1B model.
GenEval 0.605 0.601 0.569
Effect of the scaling factors in MM-RoPE. figure˜E1(b) presents the validation loss curves of varying the scaling factors for modeling temporal, height and width positional information in MM-RoPE. Two clear trends emerge: Moving from (1,1,1)(1,1,1)(2,4,4)(2,4,4)$(4,8,8)$ steadily lowers the curve, but further enlarging to $(8,16,16)$ or $(16,32,32)$ yields no additional gain since the three curves almost overlap throughout training. Therefore, a moderate scale of $(4,8,8)$ is sufficient to balance vision-language ranges and fully realize the benefit of higher-resolution RoPE, while avoiding unnecessary frequency inflation. We therefore adopt $(4,8,8)$ as the default scaling for MM-RoPE.
E.2 More Analysis
Robustness to aspect ratios. Although the aspect ratio of training data is mostly $7:4$, Tab.˜E1 indicates that Lumos-1 1B adapts well to visual generation with different aspect ratios due to the unified codebook design.
Table E2: Design comparison with other types of RoPE.
RoPE Type Compatiable with Text RoPE 3D Structure Comprehensive Frequency Allocation Strategic Scaling
M-RoPE
U-RoPE
IL-RoPE
VideoRoPE
HoPE
$M$-RoPE
Comparison with other RoPE designs. To demonstrate the superiority of MM-RoPE, we compare it with five other RoPEs, i.e., M-RoPE [80], VideoRoPE [84], U-RoPE [68], IL-RoPE [39] and HoPE [35]. We show their design comparison in Tab.˜E2. We mainly compare these methods on four aspects: 1) their compatibility with text RoPE, so that models can align with contemporary LLMs; 2) the inclusion of 3D information into RoPE to enable enhanced spatiotemporal modeling; 3) the comprehensive frequency allocation that enables more balanced local-global context modeling; 4) the strategic scaling of spatiotemporal indices (or their subsets) to account for different granularities of vision and text information. Based on the above analysis, MM-RoPE holistically incorporates the four properties.
Table E3: Performance comparison with other types of RoPE. Results are measured on GenEval, Vbench-Overall Consistency (OC), and Vbench-Imaging Quality (IQ).
RoPE Type GenEval Vbench-OC Vbench-IQ
M-RoPE 0.310 0.122 0.350
U-RoPE 0.402 0.165 0.423
IL-RoPE 0.541 0.225 0.513
VideoRoPE 0.569 0.243 0.540
HoPE 0.570 0.246 0.545
$M$-RoPE 0.591 0.249 0.559
Moreover, we conduct a quantitative comparison in Tab.˜E3, using our 1B model to perform joint training for 10k steps. Note that M-RoPE is our baseline, equal to eliminating the distributed design and the scaling design from MM-RoPE. U-RoPE improves upon M-RoPE by reallocating high-frequency components for modeling local spatial details, which is useful. IL-RoPE allocating partial channels for modeling spatial information with an interleaved design, while it only allocates partial channels for modeling texts and visual data, hurting the performance. The concurrent work, VideoRoPE, proposes several techniques for understanding tasks: low-frequency temporal allocation, diagonal layout and adjustable temporal spacing. We use VideoRoPE’s default hyperparameters. We observe that VideoRoPE slightly trails MM-RoPE in our task. We attribute this to the difference in frequency allocation and adjustable temporal spacing. Low-frequency temporal allocation enables more local modeling for spatial features and global modeling for temporal features, while it is important to maintain both local and global modeling for spatiotemporal features, which MM-RoPE performs better at. Moreover, adjustable temporal spacing increases the resolution for better modeling, and we believe that adjusting the hyperparameter in adjustable temporal spacing can increase its performance, which is a future work for VideoRoPE. HoPE improves upon VideoRoPE with a zero-frequency strategy for temporal modeling, therefore slightly outperforming VideoRoPE. It trails MM-RoPE because the frequency spectra that HoPE use for modeling spatiotemporal dependency are not as comprehensive as MM-RoPE.
Table E4: Comparison of training efficiency for text-to-image and text-to-video models.
[…]
SD-1.5 0.9B + 0.1B 781.2 0.43
SD-2.1 0.9B + 0.3B 1041.6 0.50
Dall-E 2 4.2B + 1.0B 5208.3 0.52
PixArt-$\alpha$ 0.6B + 4.7B 94.1 0.48
Lumina-mGPT 7B 4488.8 0.56
Lumos-1 1.5B 25.9 0.725
[…]
Open-Sora Plan v1.2 2.7B + 13B > 592 75.98
Lumos-1 1.5B 85.4 78.17
Comparison of training resources. We compiled a detailed comparison of training costs against popular text-to-image and text-to-video models. To ensure a fair comparison, we normalized all training times to 8xA100 GPU-Days based on their respective FP16 Tensor Core performance. The results are placed in Tab.˜E4. We can observe from the table that, compared with text-to-image diffusion models or larger-scale AR models, Lumos-1 is more compute-efficient in terms of training and obtains better performance on GenEval. Compared with Open-Sora Plan v1.2, Lumos-1 achieves similar performance, while reducing the training cost significantly. Unlike diffusion models, which rely on a powerful, external text encoder, Lumos-1 learns language understanding from scratch. Its high efficiency, even with this added learning burden, highlights the effectiveness of our discrete tokenization and MM-RoPE alignment strategy.
E.3 More Visualizations
@@
Qualitative visual comparison on text-to-image generation. We compare Lumos-1 with popular text-to-image generation methods in figure˜E2. All models are configured to their default inference settings and use the same version of detailed prompts to ensure quality generation. We can observe that: 1) Although EMU3 is significantly larger than Lumos-1, Lumos-1 generates images that reproduces fine prompt details more faithfully, such as the dappled shadows on the bench (example 1) and the correct “STOP” text and the yellow color on the road sign (example 3). EMU-3 sometimes drops or warps these details and exhibits occasional anatomical distortions, such as the distorted zebra (example 5) and malformed fingers on the woman holding an apple (example 7). 2) Stable Diffusion 3 offers slightly crisper textures, which is expected from its continuous tokenizer, but it frequently ignores descriptive constraints, such as the dappled shadows (example 1) and gentle shadows (example 4), and the incorrect dress color (example 7). By contrast, Lumos-1 meets these constraints while maintaining a convincing overall composition. These examples illustrate that Lumos-1 delivers strong vision–language alignment compared with both diffusion and autoregressive baselines, while remaining competitive in visual quality.
@@
Figure E2: Visual comparison of Lumos-1 with EMU3 and Stabel Diffusion 3 (Medium) on the text-to-image task. We place the simple version of the prompts above the images for reference. Fragments of the detailed prompt are placed in parentheses, with the critical attributes highlighted in red. All images are generated using the detailed prompts for all models.
@@
More text-to-image visualizations. figure˜E3 presents more results generated by Lumos-1 using descriptive captions. Due to the unified codebook, we can train Lumos-1 using the original resolution of the visual data, enabling generation with different aspect ratios such as $4:7$, $7:4$, and $1:1$, although we do not have a significant amount of data with aspect ratio $4:7$. Even though Lumos-1 is trained from scratch, it transforms lengthy descriptions into complex scenes that are aligned faithfully with the prompts. This can be exemplified by the second row of images, where these images clearly reflect the prompts:
@@
Figure E3: More text-to-image results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The resolutions for three rows of images are $384\times 672$, $672\times 384$ and $512\times 512$.
More image-to-video and text-to-video visualizations. figure˜E4 presents more image-to-video results and figure˜E5 presents more text-to-video results generated by Lumos-1. The prompts span a wide semantic and dynamical range, containing 1) close-up animal/human scenes (example 4, 5 in figure˜E5), 2) fast human and vehicle motion (example 1, 4 in figure˜E4 and example 1, 2 in figure˜E5), 3) fluid dynamics (example 1, 5 in figure˜E4 and example 2 in figure˜E5), and 4) complex multi-object collective motion (example 3 in figure˜E4 and example 3 in figure˜E5). The results demonstrate that Lumos-1 can 1) handle videos of different aspect ratios and 2) generate single-subject and multi-subject motion that is temporally coherent and physically plausible.
Figure E4: More image-to-video results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The results encompass videos of resolution $672\times 384$ and $768\times 320$.
Figure E5: More text-to-video results generated by Lumos-1. All images are generated using detailed prompts, but we only place short prompts here due to space limits. The resolution is $672\times 384$.
Appendix F Discussions
F.1 A Road Map to Unified Models for Understanding and Generation
Building upon our current work, a complete unified model can be achieved through a principled, staged approach:
  • Stage 1: Foundational Model Selection and Codebook Unification. The starting point is a base model with strong language capabilities. There are two primary paths: 1) Path A (Pragmatic Start): Initialize with a pre-trained VLM that already possesses strong visual and textual understanding. The main task is then to integrate our generative capabilities by extending its vocabulary with a discrete visual codebook and fine-tuning with our proposed AR-DF and MM-RoPE on generation tasks. 2) Path B (Purist, End-to-End): Start with a pure LLM. This requires developing a dual-purpose discrete visual tokenizer—one that produces tokens rich enough for both high-fidelity reconstruction (for generation) and high-level semantics (for understanding). This is a challenging but crucial research direction for true end-to-end unification.
  • Stage 2: Multi-Task Data Curation. A successful unified model requires a diverse, high-quality dataset covering all target modalities and tasks. This involves curating a mixture of: 1) Language-only Data: To maintain and enhance the model’s core reasoning and prevent catastrophic forgetting. 2) Image/Video-Text Pairs: For training both understanding (e.g., captioning, VQA) and generation (T2I, T2V) tasks. 3) Interleaved Multi-modal Documents: To teach the model to seamlessly process and generate complex sequences of text and visuals.
  • Stage 3: Joint Multi-Task and Staged Training. The final step is to train the model to be a versatile multi-task model. This should be done in stages: 1) Understanding Foundation (if using Path B): First, train the LLM on understanding tasks (e.g., VQA, captioning) to establish a strong vision-language alignment baseline. 2) Joint Generation and Understanding Training: In the main stage, train the model jointly on a mix of understanding and generation objectives. Our proposed AR-DF and MM-RoPE would serve as the core engine for all generative and understanding tasks. By sharing the same transformer backbone, the model is forced to develop a rich, shared representation space where concepts learned from understanding tasks can positively transfer to generation, and vice versa.
F.2 Limitations and Future Work Discussions
@@
F.3 Potential Societal Impacts and Safeguards
@@
Appendix G Disclosure of LLM Usage
We utilized a LLM to assist in the preparation of this manuscript. The primary role of the LLM was to improve the language, clarity, and readability of our writing. All intellectual contributions, including the research ideation, experimental design, data analysis, and the core arguments presented in this paper, are entirely our own. The LLM was used solely as a writing-enhancement tool and did not contribute to the scientific aspects of the work.